49. A metal sheet 27 cm long, 8 cm broad and 1 cm thick is melted into a cube. The difference between surface areas of two solids is ____ .

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: Explaination: [286 cm²] Hint: Surface area of the metal sheet = 2 [27 × 8 + 8 × 1 +27 × 1] = 2[216 + 35] = 502 cm² Since this sheet is melted to form a cube, therefore, Volume of cube = Volume of cuboid metal sheet ⇒ side3 = 27 × 8 × 1 cm3 ∴ side = 27×8×1−−−−−−−−√3 = 6 cm ∴ Surface area = 6 x 6² = 216 cm² Difference = 502 – 216 = 286 cm².